OpenStudy (anonymous):

how can you write a ratio of 210 and 1260

OpenStudy (karatechopper):

21:124

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

21:124 ???

OpenStudy (karatechopper):

i simpilfied it thts y

OpenStudy (anonymous):

u simplified it wrong..

OpenStudy (anonymous):

how do u solve it

OpenStudy (karatechopper):

what?!

OpenStudy (anonymous):

Assuming u want it little over big 210/1260 = 21/126 = 7/42 =1/6

OpenStudy (anonymous):

so its 1:6

OpenStudy (anonymous):

????

OpenStudy (karatechopper):

ya so hell ppl just write tht as ur answer

OpenStudy (anonymous):

1/6 is also a ratio but if u want to write it as 1:6 that's

OpenStudy (anonymous):

Ok as well.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

ur welcome.